OpenStudy (anonymous):

solve (x-2)^2 would this equal x^2-4x+4 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(x-2)^2\] \[(x-2)(x-2)\] \[x^2 -2x -2x + 4\] \[x^2 -4x + 4\]

OpenStudy (anonymous):

@satellite73 i need help with this one What is the area of a square that has a side equal to 2x+1 (A=S^2)

OpenStudy (anonymous):

\((2x+1)^2=(2x+1)(2x+1)=?\)

OpenStudy (anonymous):

i am sure you can do it, let me know what you get

OpenStudy (anonymous):

4x^2

OpenStudy (anonymous):

ok that is the first term there are some more

OpenStudy (anonymous):

1) \(2x\times 2x=4x^2\) 2) \(1\times 2x=?\)

OpenStudy (anonymous):

4x^2+2x+1?

OpenStudy (anonymous):

you missed a \(2x\) in there somewhere

OpenStudy (anonymous):

\[(2x+1)(2x+1)=2x\times 2x+1\times 2x+1\times 2x+1\times 1\]

OpenStudy (anonymous):

so its 4x^2+4x+1

OpenStudy (anonymous):

or \[4x^2+2x+2x+1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks c: